OpenStudy (anonymous):

Integral e^x/(1+e^2x) dx

OpenStudy (anonymous):

e^x = t e^xdx = dt Your integral become: dt/1+t^2

OpenStudy (anonymous):

@ncorrea92

OpenStudy (anonymous):

arctg t +C = arctg e^x +C

OpenStudy (anonymous):

Im sorry i don't understand it

OpenStudy (anonymous):

make change of variable e^x = t e^xdx = dt

OpenStudy (anonymous):

Your integral become: dt/1+t^2

OpenStudy (anonymous):

and just solve it

OpenStudy (anonymous):

(e^x)^2=e^2x

OpenStudy (turingtest):

another (probably more complicated) approach is to use the trig sub\[e^x=\tan x\]